OpenStudy (anonymous):

Solve the liner equation -7/8=-5/6z

OpenStudy (tkhunny):

What is the Leact Common Multiple of 6 and 8?

OpenStudy (anonymous):

can you help me with this please

OpenStudy (anonymous):

Cross mutiply.... so it becomes -7 * 6z = 5 * 8 therefore, -42z = -40 now the negative signs cancel out...z= 40/42 now pls solve the rest...

OpenStudy (tkhunny):

I can help you. The first step is for you to answer my question.

OpenStudy (anonymous):

what is ur question @thunny ?

OpenStudy (tkhunny):

It's up there in post #2.